OpenStudy (diogop110):

int sqrt(x+y) dx dy, y=0..1, x=1..2 ?

OpenStudy (diogop110):

\[\int\limits_{0}^{1}\int\limits_{1}^{2} \sqrt(x+y) dx dy\]

OpenStudy (diogop110):

Sorry, it's not right.

OpenStudy (zpupster):

i apoligize i had forgot the sqrt the first time